OpenStudy (anonymous):

"An archer releases an arrow from a shoulder height of 1.30 m. When the arrow hits the target 18m away, it hits point A. When the target is removed, the arrow lands 45m away. Find the maximum height of the arrow along its parabolic path." How do you find the maximum?

OpenStudy (perl):

this question was asked yesterday

OpenStudy (anonymous):

\[y - y _{o}= vt + \frac{ 1 }{ 2 }at ^{2}\] \[t = \sqrt{\frac{ 2(-y) }{ g }}\] \[t = \sqrt{\frac{ 2(-1.3m) }{ -9.8m/s ^{2} }}\]= 0.52s Now plug in the time for the x direction. \[x = vt\] \[v = \frac{ x }{ t }\] \[v = \frac{ 45m }{ 0.52 s }= 86.5m/s\] Use vx to find the angle vx cos(x) Then find the y component of v. vy sin(x) Then plug in vy back into the first equation.

OpenStudy (anonymous):

Thank you all:)

OpenStudy (perl):

you don't need to use gravity in the problem,

OpenStudy (perl):

you are given three points and you want to fit a quadratic curve to it

OpenStudy (perl):

@Jhannybean also the question is incomplete, he did not define 'point A'

OpenStudy (anonymous):

Jhannybean did not do this right. Shes using gravity in the horizontal direction. This is a projectile, the only accelerate is going on is gravity which is in the vertical direction.

OpenStudy (jhannybean):



OpenStudy (anonymous):

Looking back the question is unclear. It gave us two different ranges. The max height when the range is 18m away or the max height when the target is gone, and the range is 45m away?

OpenStudy (jhannybean):

typo.